OpenStudy (anonymous):

First estimate the answer. Then subtract to find the exact answer. 28 7/9 - 5 5/6 = 28 7/9 rounds to ? 5 5/6 rounds to ?

OpenStudy (anonymous):

\[ \left( 28 \times 9 + 7 \right)- \left( 5 \times 6 + 5 \right) = \frac{ 259 }{ 9 } - \frac{ 35 }{ 6 } = 24,88\]

OpenStudy (anonymous):

errata, 22,94